OpenStudy (anonymous):

Vertex of y= |-3x+2|

OpenStudy (anonymous):

In order to solve for the vertex, find where both sides of the absolute value equation are equal to each other. So, -3x+2=3x-2 -3x-3x=-2-2 -6x=-4 x=4/6 or x=2/3 Now, you can put x=2/3 into either equation to solve for y. I'm going to put it in the first one. y=-3x+2 y=-(3*(2/3))+2 y=-2+2 y=0 So, your vertex will be at (2/3,0). Hope that helps!